OpenStudy (tmagloire1):

If 1,200 cm2 of material is available to make a box with a square base and an open top, find the maximum volume of the box in cubic centimeters. Answer to the nearest cubic centimeter without commas. For example, if the answer is 2,000 write 2000.

OpenStudy (amistre64):

is this a calculus problem?

OpenStudy (tmagloire1):

yes calc ab.

OpenStudy (tmagloire1):

I got 45,000

OpenStudy (amistre64):

what is your equation for volume?

OpenStudy (tmagloire1):

300x - x^3/2

OpenStudy (amistre64):

1,200 = b^2 + 4bh hmm, have you done langrange multipliers?

OpenStudy (tmagloire1):

langrange multipliers? Never heard of them. this is just applied min and max problems

OpenStudy (amistre64):

V = b^2h ; A = b^2+4bh-1200 = 0 Vb = 2bh; L~Ab = L(2b+4h) Vh = b^2; L~Ah = L(4b) equating partials we have 4Lb = b^2, b=4L 2bh = 2Lb+4Lh bh-2Lh = Lb h(b-2L) = Lb h = Lb/(b-2L) = 2L ---------------- 3L^2 = 75 L=5 -------------- b=20, h=10 is what im getting ... if i did a proper langrange :)

OpenStudy (amistre64):

i tried doing a min max setup but i wasnt getting anywhere useful, so i attempted the langrange

OpenStudy (tmagloire1):

Okay that's long-range that's what we do as well. so because b=20 and h=10 the volume then we get lwh=20(10)(5)=1000?

OpenStudy (amistre64):

we could take our volume in terms of b given that b^2+4bh-1200 = 0 h = (1200-b^2)(4b)^(-1) V = b^2(1200-b^2)(4b)^(-1) or V = (1200b -b^3)/4

OpenStudy (amistre64):

no, 20*20*10

OpenStudy (tmagloire1):

Did you get the second 20 from b^2+4bh-1200 = 0

OpenStudy (amistre64):

V = (1200b -b^3)/4 V' = (1200-3b^2)/4 = 0 1200-3b^2 = 0 1200 = 3b^2 400 = b^2 ; b=20

OpenStudy (tmagloire1):

ohh ok i forgot to take the derivative of v.

OpenStudy (amistre64):

no, i got b (the base) from the workings ... the base has a side length of 20

OpenStudy (tmagloire1):

Okay, I understand now, thank you for helping me!

OpenStudy (amistre64):

youre welcome